OpenStudy (anonymous):

A quiz consists of 10 questions. Each question is marked either “right” or “wrong.” For each student in the class, the instructor keeps a record of the number of right answers the student got. In the class, the average number of right answers is 6.2 and the SD of the number of right answers is 1.4. Find the SD of the number of wrong answers.

Directrix (directrix):

Insight from @Phone_A_Friend_SS_from_California "I think the answer is 1.4. If x is the number right score than 10-x would be the number wrong score. If M is the st. dev. of right scores then 10-M is the mean of the wrong scores. (assuming here that any non-right response, including leaving answer space blank, is a wrong score. The sd. dev. formula for number right involves summing (x-M)^2. The sd. dev. formula for number wrong scores involves summing [(10-x) - (10-M)]^2 = (-x+M)^2 = (x-M)^2." @Khlara

OpenStudy (anonymous):

thanx

Directrix (directrix):

Glad the friend could help.

OpenStudy (anonymous):

There are 26 letters in the English alphabet. Letters are picked one by one at random, so that each letter has the same chance of appearing as any other letter, regardless of which letters have appeared or not appeared. a. Six letters are picked. Find the chance that the sequence that appears is RANDOM, in that order. b. Six letters are picked. Find the chance that they can be arranged to form the word RANDOM.

Directrix (directrix):

Take a look at this discussion of that problem: http://openstudy.com/study#/updates/516c4bf3e4b063fcbbbc68f6